OpenStudy (anonymous):

One angle has twice as many degrees as another angle. The sum of the degrees in both angles is 150 degrees. How many degrees are in each angle?

OpenStudy (anonymous):

Since we're talking about two angles, let the angles have x and y degrees "One angle has twice as many degrees as another angle" : x=2y (We arbitrarily assume x greater than y. As both are variables, we can do so) The sum of the degrees in both angles is 150 degrees x+y=150 Now, using x=2y, 2y+y=150 So, y=50 degree, x=2y=2*50=100 degree

OpenStudy (anonymous):

gimme medal pls